# Improved Performances and Motivation in Intelligent Tutoring Systems: Combining Machine Learning and Learner Choice

### v1

@@ 5.5 Kidlearn activities scenario
The application proposes exercises to students in the form of money games (see Figure 12). For each exercise type, one object is presented with a given tagged price, and the learner has to choose which combination of bank notes, coins or abstract tokens need to be taken from the wallet to buy the object, with various constraints depending on the exercise parameters.
@@ 5.6.1 Experiment design
According to the Randomized and Controlled Trial (RCT) methods, the experiment was designed to compare four experimental conditions corresponding to two manipulations, i.e. the algorithm conditions, and the object choice and non-choice conditions. A predefined sequence (called "Predef" ) is used as a algorithm baseline with a linear learning path. It is implemented as a series of activities in which the student must have 75% success over 4 activities of the same type to pass to the next activity type. This Predef sequence was designed by a professional in didactics of mathematics. The other condition is the ZPDES algorithm condition with personalized paths according to learning progress of student (using a parameterization of learning activities previously designed by the expert). Within each algorithm condition, we introduced the possibility to self-choose the object of the money exchange for manipulating intrinsic motivation thanks to self-decision making. Hence, four conditions were manipulated as follows: 1) a "Predef" condition with linear path and without self-choice ; 2) a "PCO" condition with linear path and with self-choice of objects; 3) a "ZPDES" condition with Learning-progress based personalizing, but without self-choice of objects ; and finally 4) a "ZCO" condition with Learning-progress based personalizing with self-choice of objects. Thanks to these 4 condition, we were able to assess the effect of Learning-progress based personalizing on learning and motivation performance as well as its possible synergistic effect with playful feature related to the self-choice of object for exercise.
@@ 5.6.2 Participants
Teachers from 11 primary schools with 2nd grade classes signed up to participate in the Kidlearn program in Nouvelle-Aquitaine, the South-West region of France. Additionally, we collected the consent from each participant and their parents. To assign each student to one of the 4 experimental conditions, randomization was conducted at the classroom level in order to avoid contamination effects. Although 414 students took the pre-intervention assessment, data from 147 students were excluded from the analysis as they did not complete all the sessions of the experimental protocol (Kidlearn session and/or post-test assessment). In addition, to ensure balance across conditions ex ante, we performed a pseudo-randomized selecting procedure at student level via a computer algorithm. In particular, children were first partitioned into strata according to the following variables: the gender (girl or boy), the child age (7 or 8 years old) and pre-assessment calculation score. Then, within each stratum, children were randomly assigned to a condition.
The final sample includes 265 children in 24 classes in 11 schools with 62 children for Predef, 59 children for PCO, 76 children for ZPDES and 68 children for ZCO condition. The background characteristics of the students in terms of demographic and school-related dimensions are summarized in table 4 in appendix.
@@ 5.6.3 Measurement toolkit
To have an assessment of the background of each participant, several measures have been collected. First, the General Profile (GP) questionnaire refers to questions related to student information such as gender, experience with technologies, his perception and habits to use money (simple manipulation or money calculation). Another set of questions concerns the habits of choosing in life-related decision making such as clothes or food choices (self-choice score): it is used to probe the student’s self-determination trait [^1] When a behaviour is self-determined, the regulatory process is choice, but when it is controlled, the regulatory process is compliance (or in some cases defiance) [80].. Second, a school-related psychological perception assessment has been carried from the use of two questionnaires,i.e., the Quality of School-Life Scale (QSLS)[81] and the Learner Empowerment Scale (LES)[82]. The QSLS evaluates the quality of school life experienced by the student (satisfaction at school, the student’s interest in academic learning, and the nature of the student-teacher interactions / students’ attitude towards the teacher), and it’s a high predictor of disengagement behaviours in school. The LES measures the learner’s empowerment with questions such as "This course will help me achieve my future goals" and "I have the qualifications to succeed in this class". The two questionnaires were combined and reworked to produce the School Profile questionnaire (SP), which contains 10 items on a 5-point Likert scale ; 5 items are related to school while the others to relationships.
The overall profile metrics aimed to establish an initial profile for each student in order to have equivalent control and experimental groups in respect of personal factors that may affect the results of our experiment such as demographic factors, technology experience, everyday self-determined behaviors, money manipulation and calculation and school experience and perceptions (see table 4 in appendix).
@@
The exploitation of the data from the interaction of the students with the Kidlearn activity is done using two kinds of indicators tracing the learning path of each student.
The first one (Fig. 3) is used to compare quantitatively the activities made by the students. Here, two scores are built, and used in result section 2.1. The former represents the activities reached by a student in the activity space, and the latter one represents the success rate over these reached activities. For a student, these scores are defined as follows:
@@
A math-test, used as a pre-post metric to measure the students money-related computational learning, has been made by a pedagogical expert. The pre- and post-test are composed of 20 items scoring from 0 to 1 (max score is 20). The three first questions are general mathematical questions about composition of numbers. The 17 other questions are related to the manipulation money, composition of numbers, addition and subtraction, and are directly related to the activities of the Kidlearn scenario. The pre-test happens at the beginning of the first session, while the post-test happens at the end of the last session. The pre- and post-test are presented on the tablet on a dedicated interface (different from the ITS one). Each item of the test evaluates the student over knowledge and skills related to money manipulation, number composition, addition or subtraction (similar to the skills and knowledge trained in the ITS). Both tests include the same items organised in the same order but the items’ wording have randomly selected values for each item and each student (with verification that no items in the post-test have the same values in their wording as the ones in the pre-test for one student).
@@
Kidlearn-related Learning experience
Two questionnaires rates learning experience from Kidlearn game, using an emotional scale and a motivation questionnaire.
The first one assessed the emotional experience or well-being related to Kidlearn practice. It consisted of self-measurements of emotional valence elicited during each Kidlearn session. This self-measurement is a simplified version of the Self-Assessment Manikin [83]. It is presented in figure 14. The student must position a cursor between "the best time of my life" (intensive and positive emotional state) and "the worst time of my life" (intensive and negative emotional state) depending on how he/she feels right now (it is scored from -50 to 50). They have to answer to this scale at the beginning, middle and end of each one of the four sessions. A summative score is computed from all inputs (used in result section 2.3), max score is 600 across the four sessions. This provides us an approximate measure of the evolution of the well-being of the children during each session.
The second measure is a motivation questionnaire from Vallerand’s questionnaire [50, 51] providing a measure of the amount of student motivation. It is based on Self-Determination Theory[52] and is commonly used to ascertain the elicitation of intrinsic and extrinsic motivation[53]. Once adapted to a public of children users playing a serious game, children had to answer a 21 items-questionnaire on their experience playing during the last session (max score of global motivation is 21).
@@ 7.1 Population characteristics for each condition
7.2 Detailed predefined sequence
Table 5 shows the $27$ successive activities for the students following the parameters defined in section 5.5.
G1.1 G1.2 G1.3 G2.1 G2.2 G2.3 G2.4 G3.1 G3.2 G4.1 G4.2 G4.3 G4.4
Ex Type M M M MM MM MM MM R R RM RM RM RM
Difficulty 1 2 3 1 1 2 2 1 2 1 1 2 2
Cents Not - - - - - - - - - - - - -
Remainder - - - - - - - - - - Int - Int
G5.1 G5.2 G5.3 G5.4 G6.5 G6.6 G6.7 G6.8 G7.1 G7.2 G7.3 G8.5 G8.6 G8.7 G8.8
Ex Type M M M M MM MM MM MM R R R RM RM RM RM
Difficulty 4 5 5 6 3 3 4 4 3 3 4 3 3 4 4
Cents Not x€x x€x x,x€ x,x€ - - - - x€x x€x x,x€ - - - -
Remainder - - - - - Int - Dec Int - Int - - Int Dec
Money Type Real Real Real Real Real Real Real Token Real Real Real Real Real Real Token
Table 5: Detailed predefined sequence

### v2

@@ 5.2 Kidlearn activities scenario
The application proposes exercises to students in the form of money games (see Fig. 11). For each exercise type, one object is presented with a given tagged price, and the learner has to choose which combination of bank notes, coins or abstract tokens need to be taken from the wallet to buy the object, with various constraints depending on the exercise parameters.
@@ 5.3.1 Experiment design
According to the Randomized and Controlled Trial (RCT) methods, the experiment was designed to compare four experimental conditions corresponding to two manipulations, i.e. the algorithm conditions, and the object choice and non-choice conditions. A predefined sequence (called "Predef" ) is used as a algorithm baseline with a linear learning path. It is implemented as a series of activities in which the student must have 75% success over 4 activities of the same type to pass to the next activity type. This Predef sequence was designed by a professional in didactics of mathematics. The other condition is the ZPDES algorithm condition with personalized paths according to learning progress of student (using a parameterization of learning activities previously designed by the expert). To estimate the necessary sample size, an a priori power analysis was conducted using G*Power 3.1.
A recent meta-analysis on studies investigating ITS in learning reported a smallest effect size at g = 0.30 [98]. Based on this effect size, with an alpha level of 0.05, the projected total sample size required for detecting a between-between interaction in a repeated measures ANOVA was approximately N = 196 for a power of 0.95, and N = 128 for a power of 0.80.
Within each algorithm condition, we introduced the possibility to self-choose the object of the money exchange for manipulating intrinsic motivation thanks to self-decision making. Hence, four conditions were manipulated as follows: 1) a "Predef" condition with linear path and without self-choice ; 2) a "PCO" condition with linear path and with self-choice of objects; 3) a "ZPDES" condition with Learning-progress based personalizing, but without self-choice of objects ; and finally 4) a "ZCO" condition with Learning-progress based personalizing with self-choice of objects. Thanks to these 4 condition, we were able to assess the effect of Learning-progress based personalizing on learning and motivation performance as well as its possible synergistic effect with playful feature related to the self-choice of object for exercise.
@@ 5.3.2 Participants
Teachers from 11 primary schools with 2nd grade classes signed up to participate in the Kidlearn program in Nouvelle-Aquitaine, the South-West region of France. Additionally, we collected the consent from each participant and their parents. To assign each student to one of the 4 experimental conditions, randomization was conducted at the classroom level in order to avoid contamination effects.
Although 414 students took the pre-intervention assessment, data from 147 students were excluded from the analysis as they did not complete all the sessions of the experimental protocol (Kidlearn session and/or post-test assessment). In addition, to ensure balance across conditions ex ante, we performed a pseudo-randomized selecting procedure at student level via a computer algorithm. In particular, children were first partitioned into strata according to the following variables: the gender (girl or boy), the child age (7 or 8 years old) and pre-assessment calculation score. Then, within each stratum, children were randomly assigned to a condition.
The final sample includes 265 children in 24 classes in 11 schools with 62 children for Predef, 59 children for PCO, 76 children for ZPDES and 68 children for ZCO condition. The background characteristics of the students in terms of demographic and school-related dimensions are summarized in table 5 in 5.4.
@@ 5.3.3 Measurement toolkit
To have an assessment of the background of each participant, several measures have been collected. First, the General Profile (GP) questionnaire (7.1.2) refers to questions related to student information such as gender, experience with technologies, his perception and habits to use money (simple manipulation or money calculation). Another set of questions concerns the habits of choosing in life-related decision making such as clothes or food choices (self-choice score): it is used to probe the student’s self-determination trait [^1] When a behaviour is self-determined, the regulatory process is choice, but when it is controlled, the regulatory process is compliance (or in some cases defiance) [99]..
Second, a school-related psychological perception assessment has been carried from the use of two questionnaires,i.e., the Quality of School-Life Scale (QSLS)[100] and the Learner Empowerment Scale (LES)[101].
The QSLS evaluates the quality of school life experienced by the student (satisfaction at school, the student’s interest in academic learning, and the nature of the student-teacher interactions / students’ attitude towards the teacher), and it’s a high predictor of disengagement behaviours in school.
The LES measures the learner’s empowerment with questions such as "This course will help me achieve my future goals" and "I have the qualifications to succeed in this class". The two questionnaires were combined and reworked to produce the School Profile (SP) questionnaire (7.1.3), which contains 10 items on a 5-point Likert scale ; 5 items are related to school while the others to relationships.
The overall profile metrics aimed to establish an initial profile for each student in order to have equivalent control and experimental groups in respect of personal factors that may affect the results of our experiment such as demographic factors, technology experience, everyday self-determined behaviors, money manipulation and calculation and school experience and perceptions (see table 5).
@@
The exploitation of the data from the interaction of the students with the Kidlearn activity is done using two kinds of indicators tracing the learning path of each student. The first one (Fig. 3) is used to compare quantitatively the activities made by the students. Here, two scores are built, and used in result section. The former represents the activities reached by a student in the activity space, and the latter one represents the success rate over these reached activities. For a student, these scores are defined as follows:
@@
A math-test, used as a pre-post metric to measure the students money-related computational learning, has been made by a pedagogical expert. The pre- and post-test are composed of 20 items scoring from 0 to 1 (max score is 20). The three first questions are general mathematical questions about composition of numbers. The 17 other questions are related to the manipulation money, composition of numbers, addition and subtraction, and are directly related to the activities of the Kidlearn scenario. The pre-test happens at the beginning of the first session, while the post-test happens at the end of the last session.
The pre- and post-test are presented on the tablet on a dedicated interface (different from the ITS one). Each item of the test evaluates the student over knowledge and skills related to money manipulation, number composition, addition or subtraction (similar to the skills and knowledge trained in the ITS). Both tests include the same items organised in the same order but the items’ wording have randomly selected values for each item and each student (with verification that no items in the post-test have the same values in their wording as the ones in the pre-test for one student).
@@
Kidlearn-related Learning experience
Two questionnaires rates learning experience from Kidlearn game, using an emotional scale and a motivation questionnaire. The first one assessed the emotional experience or well-being related to Kidlearn practice. It consisted of self-measurements of emotional valence elicited during each Kidlearn session. This self-measurement is a simplified version of the Self-Assessment Manikin [102]. It is presented in figure 13.
The student must position a cursor between "the best time of my life" (intensive and positive emotional state) and "the worst time of my life" (intensive and negative emotional state) depending on how he/she feels right now (it is scored from -50 to 50). They have to answer to this scale at the beginning, middle and end of each one of the four sessions. A summative score is computed from all inputs (used in result section),max score is 600 across the four sessions. This provides us an approximate measure of the evolution of the well-being of the children during each session.
@@
The second measure is a motivation questionnaire from Vallerand’s questionnaire [103, 104] providing a measure of the amount of student motivation. It is based on Self-Determination Theory[105] and is commonly used to ascertain the elicitation of intrinsic and extrinsic motivation[106]. Once adapted to a public of children users playing a serious game, children had to answer a 21 items-questionnaire on their experience playing during the last session (max score of global motivation is 21). This questionnaire, entitled Type of Motivation (TM), is presented in 7.1.1.
@@ 5.4 Population characteristics for each condition
6 Acknowledgements
7 Appendix
7.1 Questionnaires
7.1.1 Type of Motivation Questionnaire (TM)
The children can answer by yes or no to the items.
  1. 1.
    I played the game because I would have felt ashamed if I had not tried.
  2. 2.
    I played the game because I like succeeding at a game.
  3. 3.
    I played the game, but I do not know why.
  4. 4.
    I played the game because it makes me happy when I correctly answer a challenging activity.
  5. 5.
    I played the game to avoid doing an exercise the teacher would have given me.
  6. 6.
    I played the game because I am happy to learn how to handle money.
  7. 7.
    I played the game because it relaxes me when I play.
  8. 8.
    I played the game to be able to use the tablet/computer.
  9. 9.
    I played the game because I did what I was told to do.
  10. 10.
    I played the game because I am happy when I play.
  11. 11.
    I played the game to show that I can do things.
  12. 12.
    I played the game because I think that if I can handle money better, I can help a friend or my parents with it.
  13. 13.
    I played the game because I never get bored when I play.
  14. 14.
    I played the game because I like learning new things.
  15. 15.
    I played the game because it is a way for me to improve my calculations.
  16. 16.
    I do not know why I played the game; I feel like I was bored.
  17. 17.
    I played the game because I learn many things that interest me.
  18. 18.
    I played the game because I have always done well in games and I want to prove it.
  19. 19.
    I played the game to get a good score.
  20. 20.
    I played the game because it puts me in a good mood when I succeed.
  21. 21.
    I played the game because I am happy when I feel like I am getting better at a game.
7.1.2 General Profile (GP)
  1. 1.
    About gender:
    • (a) I am a boy
    • (b) I am a girl
  2. 2.
    About the use of money:
    • (a) I have used money and I like it
    • (b) I have used money but I do not like it
    • (c) I have never used money but I would like to learn
    • (d) I have never used money and I do not want to learn
  3. 3.
    Do you like doing calculations?
    • (a) No
    • (b) Yes
  4. 4.
    What screens do you use at home?
    • (a) A computer
    • (b) A tablet
    • (c) A smartphone
    • (d) A portable console
    • (e) A home console
  5. 5.
    How often do you use these screens?
    • (a) Never
    • (b) Only on weekends
    • (c) A few times a week
    • (d) Every day
  6. 6.
    At home, I choose the activities I do:
    • (a) Not at all
    • (b) Not often
    • (c) Sometimes yes, sometimes no
    • (d) Often
    • (e) All the time
  7. 7.
    Generally, I choose the food I eat:
    • (a) Not at all
    • (b) Not often
    • (c) Sometimes yes, sometimes no
    • (d) Often
    • (e) All the time
  8. 8.
    I choose who my friends are:
    • (a) Not at all
    • (b) Not often
    • (c) Sometimes yes, sometimes no
    • (d) Often
    • (e) All the time
  9. 9.
    I choose the clothes I wear:
    • (a) Not at all
    • (b) Not often
    • (c) Sometimes yes, sometimes no
    • (d) Often
    • (e) All the time
7.1.3 School Profile Questionnaire (SP)
  1. 1.
    I do interesting things in class:
    • (a) Never
    • (b) Almost never
    • (c) Sometimes
    • (d) Often
    • (e) All the time
  2. 2.
    At school, I get along with:
    • (a) Nobody
    • (b) One or two people
    • (c) Less than half
    • (d) More than half
    • (e) Everyone
  3. 3.
    I like my teacher:
    • (a) Not at all
    • (b) Not much
    • (c) A little
    • (d) A lot
    • (e) Really a lot
  4. 4.
    At school, I feel bored:
    • (a) All the time
    • (b) Often
    • (c) Sometimes
    • (d) Almost never
    • (e) Never
  5. 5.
    If school were a person, it would be:
    • (a) My enemy
    • (b) A stranger
    • (c) Someone I know
    • (d) A friend
    • (e) My best friend